# EvalLM: Interactive Evaluation of Large Language Model Prompts on User-Defined Criteria

### v1

@@ 1. Introduction
Based on these findings, we introduce EvalLM to facilitating prompt iterations by supporting the evaluation of outputs on user-defined and application-specific criteria (e.g., measuring “Object Familiarity” in scientific analogies for children). Instead of focusing on the low-level task of assessing generated outputs, EvalLM shifts designers’ focus to the higher-level process of refining prompts and criteria—representations of their plans and requirements. Inspired by the methodology for developing and validating psychometric scales (Robinson, 2018; Boateng et al., 2018) and techniques for LLM-based evaluations (Zheng et al., 2023; Ye et al., 2023; Liu et al., 2023a), EvalLM employs an LLM as both (1) an evaluation assistant, which evaluates outputs on the defined criteria, and (2) a criteria reviewer, which revises the defined criteria. To aid users in revising their prompts and criteria, the evaluation assistant explains its assessments, allowing the user to identify where prompt outputs fell short or to identify where the assistant’s interpretation of criteria misaligned with their own. Furthermore, the criteria reviewer analyses the user’s criteria to identify revisions that can lead to evaluations of outputs on more specific and fine-grained dimensions. Through iterations of this collaborative process, designers co-evolve their prompts and criteria, where prompts improve to satisfy criteria and criteria improve to discern the quality of prompts—ultimately leading to more polished applications.
To understand how prompt designers adopt automatic evaluations during prompt iterations, we conducted a within-subjects study (N=12) where participants improved and evaluated prompts for novel tasks proposed by recent HCI work. In the study, participants used both EvalLM and a baseline where they manually evaluated outputs—emulating designers’ current practice. Our study revealed that EvalLM helped participants “debug” their prompts by allowing them to quickly identify areas for improvement, and the evaluation assistant’s explanations served as feedback by helping participants think about how to make these improvements. As a result, we observed that participants reached satisfactory prompts more efficiently as they tested 59% fewer changes compared to when they did not have evaluation assistance. As EvalLM also facilitated criteria revision, participants felt higher satisfaction regarding the quality of their criteria—suggesting that these criteria could be valuable during human evaluations. Overall, these findings suggest that EvalLM can fill the current gap between application development and deployment by assisting designers to iterate on prompts until a stage where they have the confidence to commit resources for more robust human evaluations.
@@
EvalLM, an interactive system that aids users in revising prompts and verifying the effect of revisions by employing an LLM-based evaluation assistant to assesses outputs on user-defined criteria, and a criteria reviewer to refine these criteria to assess more specific and detailed dimensions of outputs.
@@ 2.1. Designing LLM Prompts
Although LLMs facilitate the development novel AI-based applications without the need for data collection or model training, designing satisfactory prompts can be an arduous task (Liu et al., 2023c). The specific format, phrasing, content, examples, or even the order of examples used in prompts can significantly affect performance (Lu et al., 2021; Liu et al., 2021; Rubin et al., 2021; Zhao et al., 2021; Liu et al., 2023c). However, as the space of possible natural language instructions is near infinite, designers need to test as many possibilities as possible to identify high-performing prompts (Zamfirescu-Pereira et al., 2023a; Liu et al., 2023b). To help designers identify effective prompts, researchers have proposed various tools that facilitate prompt design. For example, AI Chains (Wu et al., 2022) helps users decompose complex tasks into a chain of prompts that can be tested individually, and Kim et al. (Kim et al., 2023b) proposed a design framework for interfaces that support end-users to testing and experimentation with prompts. To automate testing, various systems (Strobelt et al., 2023; Wu et al., 2023; Mishra et al., 2023) allow users to create prompt variants and automatically compare their performance. However, while these prior approaches focused on classification tasks, evaluating performance in open-ended generative tasks can be more complex as outputs need to be examined on subjective criteria, which typically requires human effort. Building on these systems, our work proposes a human-AI collaborative system where users define subjective evaluation criteria and an LLM automatically assesses outputs on these criteria to provide insight into the performance of prompts and surface necessary improvements.
@@ 2.2. Natural Language Generation and Evaluation
Natural language generation (NLG) is the family of NLP tasks where the goal is to generate text that satisfies a communicative goal (e.g., summarize a document) while possessing several desired qualities (e.g., fluent, coherent) (Gehrmann et al., 2023; Nallapati et al., 2016; Gopalakrishnan et al., 2023; Fan et al., 2018). While recent years have shown significant progress in NLG, especially due to LLMs, a constant obstruction to progress has been the difficulty of evaluating these tasks (Khashabi et al., 2021; Dou et al., 2021; Xiao et al., 2023). Unlike classification tasks where performance is measured by comparing a predictions to a ground-truth label, generation tasks are ill-posed—i.e., multiple dissimilar outputs can be equally valid. While researchers have proposed automatic metrics that compare outputs to several ground-truth references (e.g., BLEU (Papineni et al., 2002), ROUGE (Lin, 2004)), the space of valid outputs in open-ended generative tasks can be overwhelmingly vast, making it nearly impossible to create sufficiently comprehensive references sets. Thus, human evaluations where annotators rate or rank generated text have become the golden standard in the domain (Gehrmann et al., 2023). However, the cost and effort involved in recruiting human annotators can make this type of evaluation prohibitive during earlier development stages. As an alternative, recent work (Li et al., 2023; Zheng et al., 2023; Fu et al., 2023; Liu et al., 2023a; Ye et al., 2023; Chiang and yi Lee, 2023) has employed LLMs to simulate annotators and automatically evaluate outputs on their general quality or a pre-defined set of criteria—demonstrating agreement with human evaluations on par with the level of agreement between human evaluators (Zheng et al., 2023). Inspired by these approaches, we incorporate LLM-powered simulated evaluators to support interactive evaluation of LLM prompts on user-defined criteria, and investigate how users interact with these evaluations to refine prompts.
@@ 2.3. Interactive Machine Learning Evaluation
Evaluation is fundamental to the development of machine learning models for real-world applications. Beyond assessing performance on a single metric, practitioners (e.g., developers, researchers, engineers) assess more fine-grained model behaviors to identify flaws and potential improvements (Ribeiro et al., 2020). To support fine-grained assessments, prior work has introduced various systems that allow practitioners to interactively evaluate outputs. For example, Zeno (Cabrera et al., 2023), the What-If Tool (Wexler et al., 2020), and Errudite (Wu et al., 2019) help practitioners to identify slices or subsets of data that may reveal distinct model failures. Beyond testing on existing input data, Polyjuice (Wu et al., 2021) and AdaTest (Ribeiro and Lundberg, 2022) allow practitioners to generate potentially challenging input data to test a model’s behavior. To aid practitioners to resolve issues after models are deployed, Angler (Robertson et al., 2023) combines online and offline data to help practitioners prioritize performance issues, and Deblinder (Cabrera et al., 2021) allows practitioners to collect and analyze model failure reports from crowdworkers. Our work expands on these ideas by allowing practitioners to interactively evaluate prompt outputs on fine-grained criteria with the aid of an LLM-based evaluator, which could simulate the feedback provided by end-users.
@@ 3.2.4. Evaluation to Refinements
Through the evaluations, designers identified what criteria the generated outputs failed to satisfy, and they attempted to refine their prompts to improve on these dimensions. However, most designers (P1-7) mentioned how they were unsure about how they should revise their prompts—a well-known challenge with LLM prompts (Zamfirescu-Pereira et al., 2023a). Designers mentioned how they had no alternative, but to simply test different changes and to manually evaluate outputs again to check the effect of these changes. As this involved significant effort, designers mentioned that they could struggle to verify how much a revision improved quality on specific criteria (P3-4, P6). Due to the overall complexity of the evaluation-refinement process, designers also mentioned how they would fail to record all of their prompt revisions and their effects, which prevented them from learning from previous attempts and tracking their progress (P4-5, P7).
4. Design Considerations
@@
To support prompt iteration, we must help designers to efficiently evaluate various outputs on various criteria, while also allowing them to flexibly define these criteria. Recent work (Ye et al., 2023; Liu et al., 2023a) has demonstrated that LLMs can evaluate outputs on various subjective criteria when provided with natural language descriptions of said criteria. By leveraging LLMs as evaluation assistants, we could enable designers to efficiently evaluate larger samples on diverse criteria of their own design. However, for designers to effectively use LLMs as evaluators, they must be able to define these criteria and verify that the corresponding evaluations align with their expectations. To scaffold these interrelated processes, we take inspiration from the process of developing and validating psychometric scales.
4.1. Psychometric Scales
A psychometric scale (Robinson, 2018) is a set of questions that collectively measure a variable (e.g., behavior, feeling, or action) that cannot be assessed directly. For example, the NASA task load index (NASA-TLX) questionnaire, which is frequently used in HCI, measures a person’s perceived workload (Hart and Staveland, 1988). Similar to these scales, evaluation criteria for generative tasks collectively measure the quality of outputs—a variable that cannot be directly measured. To support the construction of evaluation criteria, we adapt the three main stages in the development process of scales (Robinson, 2018; Boateng et al., 2018): item development, scale development, and scale evaluation.
4.1.1. Item Development
This stage involves designing the items or questions to include in the scale. Literature suggests combining both inductive (e.g., data analysis) and deductive (e.g., literature review) methods to design questions for psychometric scales. After designing the items, the items are frequently assessed by external judges to determine whether they are relevant and clearly defined.
4.1.2. Scale Development
Then, a scale is administered to a small set of respondents and their responses are used to assess the scale’s quality through multiple methods, including cognitive interviews and item reduction analysis. In cognitive interviews, respondents are asked to verbalize their mental process as they respond to the scale, which reveals whether items are producing the intended data and if there are any unclear items. In an item reduction analysis, responses are analyzed to determine whether there are items that assess similar variables and, thus, could be removed.
4.1.3. Scale Evaluation
Finally, scales are administered to a larger set of respondents to test their validity and reliability. Validity tests compare responses on the scale with ground-truth measurements of the variable or other similar variables. Reliability tests check whether the same respondent provides consistent responses at different points in time (i.e., test-retest reliability) and whether multiple respondents agree in their responses about the same subject (i.e., inter-rater reliability). In our work, we only take inspiration from reliability tests as designers do not have the ground-truth measurements needed for validity tests.
4.2. Design Goals
Based on the challenges identified from our interviews and insights from the literature on psychometric scales, we distill the following design goals:
@@ 4.2. Design Goals
DG2: Facilitate inspection of automatic evaluations through explanations. Similar to cognitive interviews in the development of psychometric scales, the automatic evaluator should explain and justify its evaluations so that designers can inspect whether the evaluations align with their expectations.
DG3: Allow for the definition of criteria based on output data and prior literature. As revealed by our interviews, prompt designers also employ both inductive and deductive methods when defining evaluation criteria as they envision new criteria while assessing outputs but also adapt criteria from prior work.
DG4: Review the user-defined criteria to identify potential revisions. Inspired by how scales are revised through reviews from external judges and analytical methods (e.g., item reduction), a system can review designers’ criteria to identify potential revisions, which could increase the effectiveness of subsequent evaluations.
DG5: Surface unreliable evaluations during large-scale evaluations. To gain a more comprehensive understanding of their prompt’s performance, designers may evaluate their prompt on large samples. As designers might not be able to verify each evaluation, a system can identify inconsistent or unreliable evaluations that designers can focus on checking.
@@ 5. EvalLM
Based on these design goals, we present EvalLM, an interactive system for iterating on prompts by evaluating multiple outputs on multiple user-defined criteria. While designers typically compose prompts and assess outputs to provide feedback to themselves, EvalLM transforms this into a collaborative process where a designer iteratively refines prompts and criteria based on feedback from an LLM (DG1). Specifically, the system employs an LLM as both an evaluation assistant and criteria reviewer. The evaluation assistant judges prompt outputs according to the user’s definitions of criteria and explains its evaluations, allowing the user to identify where their prompts fall short or where their criteria may be unclear (DG2). The user can define and revise criteria at any moment of time (DG3) and, to facilitate this process, the criteria reviewer analyzes the user’s criteria to recommend potential revisions that can enhance the detail and specificity of the evaluations (DG4). The main screen of EvalLM consists of three panels (Fig. 2): the generation panel (left), the data panel (middle), and the evaluation panel (right).
@@ 5.1. Interface
To illustrate the interactions and mechanisms in EvalLM, we walk through an example of an ML practitioner, Emily, who is designing prompts for a novel task of generating examples that help young children understand complex scientific phenomena.
The design of a prompt which is composed of three input boxes for the prompt’s name, system prompt and user prompt. Next to the name there are two buttons: a plus button for adding new prompts, and a right arrow button which opens the panel to browse other prompts.
The design of a criteria which is composed of two input boxes for the criteria’s name and description. The shown criteria is ’Simplicity’. Next to the name there is a button that shows a line splitting into two, and underneath the criteria there are two additional criteria: ’Language Simplicity’ and ’Relatability’.
@@ 5.1.1. Composing Prompts
In the generation panel, the user writes the overall instruction for their task (Fig. 2A) and designs two prompt templates (Fig. 2B). EvalLM is designed for comparing prompts as this enables designers to compare the performance of different prompt variations, or to compare prompts before and after specific edits (DG6). Furthermore, prior work has found that it is easier for both humans (Gehrmann et al., 2023; Li et al., 2019) and LLMs (Zheng et al., 2023; Bai et al., 2023) to compare model outputs than to rate a single output. For each prompt template, the user can compose the system and user prompt (Fig. 4D-E). In either of these, the user can add the tokens {{instruction}} and {{input}}, which are replaced with the instruction and the content of an input sample when the prompts are used to generate outputs. By composing the overall instruction separately, users can re-use the same base instruction across prompt templates, and test the effect of adding more information to a prompt or changing its format. To test and compare different prompt ideas, the user can create more prompts (Fig. 4B), name them, and switch between them as they desire (Fig. 4C).
In the Instruction, Emily describes her task: writing an example to help explain a piece of scientific information to a young child. To test the effect of different prompting “tricks”, Emily first creates a basic prompt that follows a simple form-like format, and creates another prompt with sub-headers, additional context, and a system prompt that instructs the LLM to act as a teacher.
A data row in EvalLM depicts, from left to right, an input sample, the output from the first prompt, the output from the second prompt, and a table of evaluations. The table of evaluations shows rows for each defined criteria and three circles where the filled in circle represents which prompt won. The explanation for one criteria is shown below the data row and parts of the output are highlighted with teh same color as the criteria.

### v2

@@ 1. Introduction
Based on these findings, we introduce EvalLM to facilitate prompt iterations by supporting the evaluation of outputs on user-defined and application-specific criteria (e.g., measuring “Object Familiarity” in scientific analogies for children). Instead of focusing on the low-level task of assessing generated outputs, EvalLM shifts designers’ focus to the higher-level process of refining prompts and criteria—representations of their plans and requirements. Inspired by recent techniques for LLM-based evaluations (Zheng et al., 2023; Ye et al., 2023; Liu et al., 2023a), EvalLM employs an LLM as both (1) an evaluation assistant, which evaluates outputs on the defined criteria, and (2) a criteria reviewer, which revises the defined criteria. To aid users in revising their prompts and criteria, the evaluation assistant explains its assessments, allowing the user to identify where prompt outputs fell short or to identify where the assistant’s interpretation of criteria misaligned with their own. Furthermore, the criteria reviewer analyses the user’s criteria to identify revisions that can lead to evaluations of outputs on more specific and fine-grained dimensions. Through iterations of this collaborative process, designers co-evolve their prompts and criteria, where prompts improve to satisfy criteria and criteria improve to discern the quality of prompts—ultimately leading to more polished applications.
To understand how prompt designers adopt automatic evaluations during prompt iterations, we conducted a within-subjects study (N=12) where participants improved and evaluated prompts for novel tasks proposed by recent HCI work. In the study, participants used both EvalLM and a baseline where they manually evaluated outputs—emulating designers’ current practice. Our study revealed that EvalLM helped participants “debug” their prompts by allowing them to quickly identify areas for improvement, and the evaluation assistant’s explanations served as feedback by helping participants think about how to make these improvements. As a result, we observed that participants reached satisfactory prompts more efficiently as they tested 59% fewer changes than when they did not have evaluation assistance. As EvalLM also facilitated criteria revision, participants felt higher satisfaction regarding the quality of their criteria—suggesting that these criteria could be valuable during human evaluations. Overall, these findings suggest that EvalLM can fill the current gap between application development and deployment by assisting designers to iterate on prompts until a stage where they have the confidence to commit resources for more robust human evaluations.
@@
EvalLM, an interactive system that aids users in revising prompts and verifying the effect of revisions by employing an LLM-based evaluation assistant to assess outputs on user-defined criteria, and a criteria reviewer to refine these criteria to assess more specific and detailed dimensions of outputs.
@@ 2.1. Designing LLM Prompts
Although LLMs facilitate the development of novel AI-based applications without the need for data collection or model training, designing satisfactory prompts can be an arduous task (Liu et al., 2023c). The specific format, phrasing, content, examples, or even the order of examples used in prompts can significantly affect performance (Lu et al., 2021; Liu et al., 2021; Rubin et al., 2021; Zhao et al., 2021; Liu et al., 2023c). However, as the space of possible natural language instructions is near infinite, designers need to test as many possibilities as possible to identify high-performing prompts (Zamfirescu-Pereira et al., 2023a; Liu et al., 2023b). To help designers identify effective prompts, researchers have proposed various tools that facilitate prompt design. For example, AI Chains (Wu et al., 2022) helps users decompose complex tasks into a chain of prompts that can be tested individually, and Kim et al. (Kim et al., 2023b) proposed a design framework for interfaces that support end-users to testing and experimentation with prompts. To automate testing, various systems (Strobelt et al., 2023; Wu et al., 2023; Mishra et al., 2023) allow users to create prompt variants and automatically compare their performance. However, while these prior approaches focused on classification tasks, evaluating performance in open-ended generative tasks can be more complex as outputs need to be examined on subjective criteria, which typically requires human effort. Building on these systems, our work proposes a human-AI collaborative system where users define subjective evaluation criteria and an LLM automatically assesses outputs on these criteria to provide insight into the performance of prompts and surface necessary improvements.
@@ 2.2. Natural Language Generation and Evaluation
Natural language generation (NLG) is the family of NLP tasks where the goal is to generate text that satisfies a communicative goal (e.g., summarize a document) while possessing several desired qualities (e.g., fluent, coherent) (Gehrmann et al., 2023; Nallapati et al., 2016; Gopalakrishnan et al., 2023; Fan et al., 2018). While recent years have brought significant progress in NLG, especially due to LLMs, a constant obstruction to progress has been the difficulty of evaluating these tasks (Khashabi et al., 2021; Dou et al., 2021; Xiao et al., 2023). Unlike classification tasks where performance is measured by comparing a prediction to a ground-truth label, generation tasks are ill-posed—i.e., multiple dissimilar outputs can be equally valid. While researchers have proposed automatic metrics that compare outputs to several ground-truth references (e.g., BLEU (Papineni et al., 2002), ROUGE (Lin, 2004)), the space of valid outputs in open-ended generative tasks can be overwhelmingly vast, making it nearly impossible to create sufficiently comprehensive references sets. Thus, human evaluations where annotators rate or rank generated text have become the golden standard (Gehrmann et al., 2023). However, the cost and effort involved in recruiting human annotators can make this type of evaluation prohibitive during early development stages. As an alternative, recent work (Li et al., 2023; Zheng et al., 2023; Fu et al., 2023; Liu et al., 2023a; Ye et al., 2023; Chiang and yi Lee, 2023) has employed LLMs to simulate annotators and automatically evaluate outputs on their overall quality or a pre-defined set of criteria—demonstrating agreement with human evaluations on par with the level of agreement between human evaluators (Zheng et al., 2023). Inspired by these approaches, we incorporate LLM-powered simulated evaluators to support interactive evaluation of LLM prompts on user-defined criteria and investigate how users interact with these evaluations to refine prompts.
@@ 2.3. Interactive Machine Learning Evaluation
Evaluation is fundamental to the development of machine learning models for real-world applications. Beyond assessing performance on a single metric, practitioners (e.g., developers, researchers, engineers) assess more fine-grained model behaviors to identify flaws and potential improvements (Ribeiro et al., 2020). To support fine-grained assessments, prior work has introduced various systems that allow practitioners to interactively evaluate outputs. For example, Zeno (Cabrera et al., 2023), the What-If Tool (Wexler et al., 2020), and Errudite (Wu et al., 2019) help practitioners to identify slices or subsets of data that may reveal distinct model failures. Beyond testing on existing input data, Polyjuice (Wu et al., 2021) and AdaTest (Ribeiro and Lundberg, 2022) allow practitioners to generate potentially challenging input data to test a model’s behavior. To aid practitioners in resolving issues after models are deployed, Angler (Robertson et al., 2023) combines online and offline data to help practitioners prioritize performance issues, and Deblinder (Cabrera et al., 2021) allows practitioners to collect and analyze model failure reports from crowdworkers. Our work expands on these ideas by allowing practitioners to interactively evaluate prompt outputs on fine-grained criteria with the aid of an LLM-based evaluator, which could simulate the feedback provided by end-users.
@@ 3.2.4. Evaluation to Refinements
Through the evaluations, designers identified what criteria the generated outputs failed to satisfy, and they attempted to refine their prompts to improve on these dimensions. However, most designers (P1-7) mentioned how they were unsure about how they should revise their prompts—the well-known challenge of prompt engineering (Zamfirescu-Pereira et al., 2023a). Designers mentioned how they had no alternative, but to simply test different changes and to manually evaluate outputs again to check the effect of these changes. As this involved significant effort, designers mentioned that they could struggle to verify how much a revision improved quality on specific criteria (P3-4, P6). Due to the overall complexity of the evaluation-refinement process, designers also mentioned how they would fail to record all of their prompt revisions and their effects, which prevented them from learning from previous attempts and tracking their progress (P4-5, P7).
4. Design Goals
@@
To support prompt iteration, we must support efficient evaluation of outputs on designers’ own criteria. Recent work (Ye et al., 2023; Liu et al., 2023a) showed that LLMs can evaluate text on diverse subjective criteria—revealing their potential as evaluation assistants. However, for designers to effectively use LLMs as evaluators, they must be able to define their own criteria and verify that subsequent evaluations align with their expectations. To scaffold these interrelated processes, we distill insights from our interviews into design goals for an LLM-based prompt iteration and evaluation system. We additionally take high-level inspiration from the process for developing psychometric scales (Robinson, 2018)—question sets that collectively measure a variable (e.g., behavior, feeling) that cannot be assessed directly (e.g., NASA-TLX (Hart and Staveland, 1988))—as we notice parallels between scale questions and evaluation criteria.
@@
DG2: Facilitate inspection of automatic evaluations through explanations. Similar to how cognitive interviews can reveal incorrect or unclear questions in scales by asking respondents to verbalize their thoughts (Boateng et al., 2018), the automatic evaluator should explain and justify its evaluations so that designers can inspect whether the evaluations align with their expectations.
DG3: Allow for the definition of criteria based on output data and prior literature. As revealed by our interviews, prompt designers envision new criteria by assessing outputs and also by referring to prior work—resembling the inductive and deductive methods for defining psychometric scale questions (Boateng et al., 2018).
DG4: Review the user-defined criteria to identify potential revisions. Inspired by how scales are revised through reviews from external judges (Robinson, 2018; Boateng et al., 2018), a system should aid designers in reviewing criteria to identify potential revisions, which could increase the effectiveness of subsequent evaluations.
DG5: Surface unreliable evaluations during large-scale evaluations. While designers need to evaluate larger samples to comprehensively assess performance, they may be unable to verify all these evaluations. As an alternative, we take inspiration from reliability tests (e.g., inter-rater, test-retest) for psychometric scales (Boateng et al., 2018; Robinson, 2018) and suggest that an evaluation system should surface less reliable evaluations for designers to verify.
@@ 5. EvalLM
Based on these design goals, we present EvalLM (Fig. 2), an interactive system for iterating on prompts by evaluating multiple outputs on multiple user-defined criteria. While designers typically compose prompts and assess outputs to provide feedback to themselves, EvalLM transforms this into a collaborative process where a designer iteratively refines prompts and criteria based on feedback from an LLM (DG1). Specifically, the system employs an LLM as both an evaluation assistant and criteria reviewer. The evaluation assistant judges prompt outputs based on the user’s definitions of criteria and explains its evaluations, which can reveal where the prompts fall short or where criteria may be unclear (DG2). The user can define and revise criteria at any time (DG3) and, to facilitate this, the criteria reviewer recommends potential revisions that can enhance the detail and specificity of evaluations (DG4).
@@ 5.1. Interface
To illustrate the interactions in EvalLM, we walk through an example of an ML practitioner, Emily, who is designing prompts for a novel task of generating examples that help young children understand complex scientific phenomena.
The design of a prompt is composed of three input boxes for the prompt’s name, system prompt and user prompt. Next to the name there are two buttons: a plus button for adding new prompts, and a right arrow button which opens the panel to browse other prompts.
The design of a criterion which is composed of two input boxes for the criterion’s name and description. The shown criteria is ’Simplicity’. Next to the name there is a button that shows a line splitting into two, and underneath the criteria there are two additional criteria: ’Language Simplicity’ and ’Relatability’.
@@ 5.1.1. Composing Prompts
In the generation panel, the user writes the overall instruction for their task (Fig. 2A) and designs two prompt templates (Fig. 2B). EvalLM is designed for comparing prompts as this enables designers to compare the performance of different prompt variations, or to compare prompts before and after specific edits (DG6). Furthermore, prior work has found that it is easier for both humans (Gehrmann et al., 2023; Li et al., 2019) and LLMs (Zheng et al., 2023; Bai et al., 2023) to compare model outputs than to rate a single output. For each prompt template, the user can compose the system and user prompt (Fig. 3D-E) using the tokens {{instruction}} and {{input}}, which are replaced with the instruction and content of an input sample when generating outputs. By composing the task instruction separately, users can reuse the same base instruction across prompt templates. To test and compare different prompt ideas, the user can create more prompts (Fig. 3B), name them, and switch between them as they desire (Fig. 3C).
Emily describes her task in the Instruction field and, to test the effect of different prompting “tricks”, she creates two prompts: a basic prompt with a simple form-like format and a prompt with sub-headers, additional context, and a system prompt that instructs the LLM to act as a teacher.
A data row in EvalLM depicts, from left to right, an input sample, the output from the first prompt, the output from the second prompt, and a table of evaluations. The table of evaluations shows rows for each defined criteria and three circles where the filled in circle represents which prompt won. The explanation for one criteria is shown below the data row and parts of the output are highlighted with the same color as the criteria.